### v2

# BRepFormer: Transformer-Based B-rep Geometric Feature Recognition

Abstract.
Recognizing geometric features on B-rep models is a cornerstone technique for multimedia content-based retrieval and has been widely applied in intelligent manufacturing. However, previous research often merely focused on Machining Feature Recognition (MFR), falling short in effectively capturing the intricate topological and geometric characteristics of complex geometry features. In this paper, we propose BRepFormer, a novel transformer-based model to recognize both machining feature and complex CAD models’ features. BRepFormer  encodes and fuses the geometric and topological features of the models. Afterwards, BRepFormer  utilizes a transformer architecture for feature propagation and a recognition head to identify geometry features. During each iteration of the transformer, we incorporate a bias that combines edge features and topology features to reinforce geometric constraints on each face. In addition, we also proposed a dataset named Complex B-rep Feature Dataset (CBF), comprising 20,000 B-rep models. By covering more complex B-rep models, it is better aligned with industrial applications. The experimental results demonstrate that BRepFormer  achieves state-of-the-art accuracy on the MFInstSeg, MFTRCAD, and our CBF datasets.
CAD, Geometric Feature Recognition, Boundary Representation (B-rep), Transformer
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; ccs: Computing methodologies Computer vision problemsccs: Applied computing Computer-aided manufacturingccs: Computing methodologies Shape modeling
1. Introduction
Geometric feature recognition serves as a critical link between Computer-Aided Design (CAD) and Computer-Aided Manufacturing (CAM). It is a cornerstone technique for multimedia content-based retrieval and plays a key role in automating manufacturing processes, improving efficiency, and reducing human errors. While traditional rule-based geometric feature recognition methods are widely used in industry, they are labor-intensive and struggle to adapt to complex geometric variations and topological changes (Henderson and Anderson, 1984; Joshi and Chang, 1988; Vandenbrande and Requicha, 1993). To address these limitations, learning-based approaches have been introduced, often converting CAD models into intermediate representations such as point clouds (Lei et al., 2022), voxels (Ning et al., 2023), or images (Shi et al., 2020b). However, these transformations lead to key topological and geometric information loss, increased computational costs, and reduced recognition accuracy. Although recent deep learning models that directly process boundary representation (B-rep) data have shown promise in preserving geometric and topological details, as well as improving recognition accuracy (Cao et al., 2020), challenges remain in handling highly complex topologies and diverse manufacturing processes, limiting their practical effectiveness.
We propose BRepFormer, a transformer-based geometric feature recognition network that leverages a transformer architecture to effectively capture and process B-rep features. Unlike previous approaches that suffer from information loss and limited topological awareness (Ma et al., 2019), BRepFormer directly operates on the boundary representation (B-rep), ensuring high-fidelity feature extraction. Specifically, our model extracts both the geometric and topological features of the CAD model from multiple perspectives. During the feature encoding stage, geometric edge features and topological features are processed by separate encoders and then fused to form an attention bias, which serves as a constraint input into the transformer module. Meanwhile, the extracted geometric face features are encoded into tokens together with a virtual face. These tokens serve as carriers of information that are fed into the transformer module, facilitating deep interaction and information fusion among the features. Finally, the recognition head accepts the output from the transformer module and fuses the global and local features within it, achieving high-precision recognition of geometric features. By integrating these components, BRepFormer effectively propagates information across the CAD model structure, improving geometric feature recognition accuracy while preserving critical geometric and topological relationships.
We conducted experiments on the public MFInstSeg (Wu et al., 2024) and MFTRCAD (Xia et al., 2024) datasets, as well as our CBF dataset. The results demonstrate that BRepFormer achieves state-of-the-art recognition accuracy on these three datasets. Notably, BRepFormer excels in recognizing complex geometric structures and diverse machining features. On the MFTRCAD (Xia et al., 2024) dataset, it achieved a 93.16% recognition accuracy, surpassing the previous best method by 3.28 percentage points. To further validate the effectiveness of our approach, we performed detailed ablation studies, analyzing the impact of each model component on overall performance. The ablation study results indicate that each feature component contributes to an improvement in the model’s accuracy.
In summary, our proposed network, BRepFormer, effectively leverages the inherent information within the B-rep structure and achieves high-precision geometric feature recognition performance. The key contributions of this work are as follows:
  • We develop a method (Sec. 3.2 and 3.3) that effectively extracts and encodes both topological and geometric features from CAD models, enabling a more informative and structured representation for machine learning models.
  • We introduce a novel transformer-based architecture (Sec. 3.4) that enforces edge and face constraints during information propagation, significantly improving feature recognition accuracy.
  • We introduce a CAD dataset (Sec. 4) that is more aligned with industrial applications, offering a more complex collection of CAD models to contribute to the geometric feature recognition.
  • Our BRepFormer  model achieves state-of-the-art accuracy on the experimental datasets, demonstrating the superiority of our approach.
Figure 1. Our model consists of four main components: (1) Feature Extractor, which extracts topological and geometrical features from the B-rep model; (2) Feature Encoder, where edge and topological features are combined to create an attention bias and face features are augmented with a virtual face before being input into the transformer block; (3) Transformer Block, which processes the encoded features using grouped query attention and attention bias; and (4) Recognition Head, where the output features are fused and passed through a classification head to obtain the recognition results.
2. Related Work
2.1. Rule-Based Geometric Feature Recognition
Rule-based geometric feature recognition methods identify geometric features in CAD models by applying predefined rules based on topological and geometric information. Early studies by Henderson (Henderson, 1984), Donaldson (Donaldson and Corney, 1993), and Chan (Chan and Case, 1994) introduced rule-based approaches that define boundary patterns and use expert systems for feature recognition. While intuitive and easy to implement, these methods struggle with flexibility due to the inherent limitations of predefined rules. Encoding all machining knowledge into a rule-based system is challenging, making maintenance cumbersome and leading to poor adaptability, particularly in handling complex intersecting features.
Volumetric decomposition is another well-established rule-based approach for geometric feature recognition. This method decomposes the material to be removed from a CAD model into intermediate volumes and reconstructs geometric features based on predefined rules. Woo et al. (Woo, 1982) introduced the Alternating Sum of Volumes (ASV) decomposition method, representing solids as a hierarchical structure of convex bodies through union and difference operations. Requicha et al. (Vandenbrande and Requicha, 1993) further developed an automatic feature recognition approach that decomposes machinable volumes into manufacturable features, addressing feature interactions using generation-testing strategies and computational geometry techniques. Wilde et al. (Kim and Wilde, 1992) refined the ASV method by introducing partitioning techniques, resolving convergence issues, and enhancing its applicability in feature recognition.
Graph-based methods have gained attention recently due to their strong alignment with the boundary representation (B-rep) structure of CAD models. These methods construct an Attribute Adjacency Graph (AAG) to capture the topological relationships and geometric attributes of faces and edges. Geometric features can then be identified by applying graph-matching techniques to detect subgraph patterns within the AAG. Joshi et al. (Joshi and Chang, 1988) first applied the AAG for topological and geometric information matching. Shah et al. (Gao and Shah, 1998) enhanced this approach by integrating hint-based feature recognition, introducing minimal condition subgraphs to improve the handling of feature interactions.
Hint-based methods (Han and Requicha, 1997; Li et al., 2015) offer a rule-driven approach for identifying complex intersecting features. They extract topological and geometric patterns, along with heuristic ”hints” derived from residuals left by feature intersections. These hints guide the reconstruction of incomplete feature information through reasoning. While this method improves the recognition of intersecting features and aids subsequent process planning, defining comprehensive and precise hints, along with robust reasoning rules, remains challenging. This complexity makes the hint-based method challenging to achieve fully automated and high-precision feature recognition in practical applications.
2.2. Deep Learning-Based Geometric Feature Recognition
Various deep learning-based methods (Lei et al., 2022; Yao et al., 2023; Xu et al., 2024; Huang et al., 2024, 2025) have been developed for 3D structural representation, each addressing different challenges. Point-cloud-based methods leverage neural networks to extract features but often suffer from information loss. MFPointNet (Lei et al., 2022) employs selective downsampling layers for feature recognition, while Yao et al. (Yao et al., 2023) proposed a hierarchical neural network to improve the recognition of complex overlapping features. Shi et al. (Shi et al., 2020a) introduced a multi-sectional view (MSV) representation and MsvNet, enriching 3D model representation by incorporating multi-view features.
Voxel-based approaches use 3D convolutional neural networks (CNNs) to process CAD models but face resolution-related information loss. FeatureNet (Zhang et al., 2018) applies 3D CNNs for feature recognition, while Peddireddy et al. (Peddireddy et al., 2020, 2021) refined voxelization techniques to predict machining processes like milling and turning. Despite these improvements, voxelization inherently reduces geometric fidelity, particularly at low resolutions. Mesh-based approaches seek to retain geometric details for improved recognition. Jia et al. (Jia et al., 2023) proposed an innovative method that combines the original MeshCNN with Faster RCNN, forming a geometric feature recognition scheme based on Mesh Faster RCNN. This approach enhances the accuracy of geometric feature detection while preserving the mesh geometry. However, the high memory demand when processing high-resolution data limits its application in large-scale scenarios.
In the CAD industry, B-rep (boundary representation) is the dominant format for 3D models, making graph-based representations particularly effective for geometric feature recognition. Graph neural networks (GNNs) have been widely applied due to their structural similarity with B-rep. Cao et al. (Cao et al., 2020) pioneered the transformation of B-rep models into graph-structured representations for learning. Colligan et al. (Colligan et al., 2022) introduced Hierarchical CADNet, a novel approach that leverages a two-level graph representation to improve recognition accuracy. Specifically, it utilizes the Face Adjacency Graph (FAG) to capture topological information and mesh patches to represent geometric details. Jayaraman et al. (Jayaraman et al., 2021) proposed UV-Net, which encodes surfaces and curves with CNNs and utilizes graph convolutional networks (GCNs) for feature learning.
Recent advancements further refine B-rep-based learning methods. Lambourne et al. (Lambourne et al., 2021) developed BRepNet, which defines convolution kernels for directed coedges, improving pattern detection. Lee et al. (Lee et al., 2023) introduced BRepGAT, incorporating graph attention networks (GATs) for precise feature segmentation. Wu et al. (Wu et al., 2024) proposed AAGNet, a multi-task GNN that simultaneously performs semantic, instance, and base segmentation using the geometric attribute adjacency graph (gAAG). Xia et al. (Xia et al., 2024) developed MFTReNet, which learns semantic segmentation, instance grouping, and topological relationship prediction directly from B-rep data. Despite these advancements, the generation of large-scale datasets with detailed topological labels requires substantial annotation efforts, thereby increasing the cost of data preparation.
3. Method
3.1. Overview
We propose a novel approach for CAD geometric feature recognition based on a transformer architecture as show in Figure 1, consisting of four different parts : 1) Feature Extractor Module considers the topological and geometric features of the model from multiple perspectives; 2)  Feature Encoder Module encodes the initial input features into a format that is friendly for the network, thereby generating the main integrated features; 3) Transformer Block Module further extracts features using a transformer structure, with a designed attention bias constraint; 4) Recognition Head Module fuses the features and classifiers the faces of B-rep data.
3.2. Feature Extraction Module
The Feature Extraction Module extracts feature for both the geometry and topology of the B-rep model, focusing on faces and edges.
3.2.1. Topological Feature Extraction
For topology, we extract four different features, including three face features (Face Shortest Distance, Face Angular Distance and Face Centroid Distance) and one feature for edges (Shortest Edge Path).
Face Shortest Distance. To fully capture the direct and indirect spatial relationships between any two faces in the B-rep model, we employed the Dijkstra algorithm to compute the shortest path length between all pairs of faces. This method quantifies the topological distance between faces as the number of intervening faces along the connecting path. Based on these results, we constructed an extended adjacency matrix $M_{d}\in R^{N\times N}$, where each element $m_{d}(f_{i},f_{j})$ represents the shortest distance from any face $f_{i}$ to another face $f_{j}$. This matrix captures not only direct connections between faces but also indirect connections via other faces, thereby reflecting complex connectivity patterns.
Face Angular Distance. To more accurately describe the relative position between two faces, our approach also extracts the dihedral angle between any two faces. When two faces share a common edge, the dihedral angle can explicitly indicate the geometric relationship formed by the two faces, whether it is concave, convex, or a smooth transition. For two non-adjacent faces, we use their normal vectors to calculate the angle between them. This method reflects the relative direction of the two faces in three-dimensional space, regardless of the series of intermediate faces through which they are connected. At the same time, our approach also constructs a matrix $M_{a}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{a}(f_{i},f_{j})$ represents the angle between a face $f_{i}$ and another face $f_{j}$.
Face Centroid Distance. The centroid distance between faces, as another key topological feature, can be used to quantify their spatial separation by measuring the Euclidean distance between the centroids of two faces. To eliminate the influence of the model’s scale, we normalize this distance by the diagonal length of the bounding box of the entire solid model, obtaining a relative distance indicator. This normalized centroid distance can reflect the similarity between CAD models of different sizes and proportions. Similarly, we construct a matrix $M_{c}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{c}(f_{i},f_{j})$ represents the Euclidean distance between the centroids of face $f_{i}$ and face $f_{j}$.
Shortest Edge Path. Considering the key role of edges in defining global topological features, our approach particularly focuses on the shortest edge path between any two faces. Edges are not only the basic elements connecting different faces but also carry rich information about the model’s internal connectivity and surface continuity. Therefore, for any two face $f_{i}$ and $f_{j}$ in the B-rep model, we not only calculate the shortest distance between them but also record all the edge chains $\left\{e_{ik},e_{kl},\ldots,e_{mj}\right\}$ that make up this shortest path. To effectively represent this edge path information, we introduce a three-dimensional matrix $M_{e}\in\mathbb{R}^{N\times N\times\text{MaxDistance}}$ to store edge path data. The parameter MaxDistance is employed to specify the upper limit of the distance between any two surfaces within a single model. For any pair of surfaces $\left(f_{i},f_{j}\right)$, in a given model, when the shortest path distance between them is less than MaxDistance, the elements in the edge path matrix that exceed the actual shortest path range will be initialized to -1 as an indicator of invalid values.
3.2.2. Geometric Feature Extraction
Our approach further conducts geometric feature extraction from two aspects: the UV domain and geometric attributes, as detailed below.
For the extraction of UV domain features from CAD models, our approach is inspired by UV-Net (Jayaraman et al., 2021). We sample and discretize the parametric surfaces and parametric curves surfaces in B-rep into regular 2D and 1D point grids with a uniform step size, as shown in Figure 2. For the grid feature representation of each face, the 3D coordinates, 3D normal vectors of the points, and an additional dimension indicating visibility are included. For the grid feature representation of each edge, the 3D coordinates and 3D tangent vectors of each point are also included. Moreover, to enhance the information representation, we add the normal vectors of the two adjacent faces of the edge to the edge grid points. Compared with discrete representation methods that may be insufficient in accurately describing complex surfaces and curves (such as voxel (Yan et al., 2023) grids and traditional meshes (Feng et al., 2019) ), our use of UV grids can capture precise geometric information and achieve a friendly input representation for neural networks.
Furthermore, our method also extracts the geometric attributes inherent in the solid entities of CAD models as show in Figure 2. To characterize the geometric attribute features of surfaces, we extract the following information: surface type (e.g., plane, conical surface, cylindrical surface, etc.), area, centroid coordinates, and whether it is a rational B-spline surface. For the geometric attribute features of curved edges, we extract their type (e.g., circular, closed curve, elliptical, straight line, etc.), length, and convexity (i.e., concave, convex, or smooth transition). These geometric attributes of faces and edges can be directly obtained from the original B-rep structure and are encoded separately. By integrating the aforementioned geometric attributes extracted from the UV domain, we obtain a geometric input representation of the entire CAD model, which provides strong support for subsequent downstream tasks.
Figure 2. The upper part of the figure shows the details of geometric UV domain sampling, while the lower part shows the details of geometric attribute sampling.
3.3. Feature Encoder Module
Feature encoder module further encodes all features extracted above, and output an attention bias and faces features, as the input for the following module.
3.3.1. Topological Feature Encoder.
For the above-mentioned three topological relation matrices $M_{d},M_{a},M_{c}\in\mathbb{R}^{N\times N}$ between faces, we applied a unified processing pipeline to these matrices. Specifically, each matrix is first transformed through a linear layer of the same dimension, followed by sequential processing through a normalization layer and a ReLU activation function for feature encoding, which is formulated as:
(1) $M_{i}^{\prime}=\text{ReLU}(\text{Norm}(\text{Linear}((M_{i})))$
here, $M_{i}$ denotes the three topological matrices $M_{d},M_{a},\text{and}\ M_{c}$.
Based on these three parts, we obtain the face bias input into the transformer module:
(2) $B_{\text{Face}}=\text{Add}({M_{d}^{\prime}},{M_{a}^{\prime}},{M_{c}^{\prime}})$
within the framework of the edge path matrix, we have considered the impact of edge weights on faces. For the shortest path between any two faces in a B-rep model, the influence of distant edges on the initial face diminishes progressively as the position along the path advances from the starting face. Based on these considerations, our method integrates the edge path matrix with edge features to model the weight influence of edges on faces along the path. The formula is expressed as follows:
(3) $B_{Edge}=\ {(M}_{e}\ \otimes H_{edge})$
here, $M_{e}$ stands for the edge path matrix, and $H_{edge}$ indicates the extracted and encoded edge features.
Finally, the sum of the obtained $B_{Edge}$ and $B_{Face}$ yields the attention bias that we input into the transformer module as shown in Figure 1.
3.3.2. Geometric Feature Encoder
For the encoding of UV domain sampling features in CAD geometric features, BRepFormer  encodes face feature as $f_{geo}\in\mathbb{R}^{N_{f}\times N_{u}\times N_{v}\times 7}$ , and edge feature as $e_{\text{geo}}\in\mathbb{R}^{N_{e}\times N_{u}\times 12}$. Here, $N_{f}$ and $N_{v}$ indicate the number of sampling points for faces and edges based on the UV grid parameters respectively. $N_{u}$ and $N_{v}$ represent the number of sampling points along the u-axis and v-axis. The aforementioned geometric features are all fed into their respective encoders. The face encoder consists of three 2D CNN layers, an adaptive average pooling layer, and a fully connected layer, which encode the face features into a feature dimension of 128. Similarly, the edge encoder has a structure akin to the face encoder but uses 1D CNN layers for preliminary encoding. The encoded information of geometric features is shown in Table 1.
Table 1. Geometric Feature Dimensions in UV Domain
Element Feature Dimension
Face Coordinates 3D
Normal vectors 3D
Visibility 1D
Edge Coordinates 3D
Tangent 3D
Normal vectors to neighboring surfaces 6D
We employed Multilayer Perceptrons (MLP) to encode the geometric attribute features of faces and edges. The encoded face attributes include a 9-dimensional one-hot vector $f_{\text{type}}\in\mathbb{R}^{9}$ to represent surface types (e.g., plane, conical surface, cylindrical surface, etc.), a vector $f_{\text{area}}\in\mathbb{R}^{1}$ to identify the surface area, a vector $f_{\text{cen}}\in\mathbb{R}^{3}$ to identify the centroid coordinates of the face, and a vector $f_{\text{rat}}\in\mathbb{R}^{1}$ to identify whether it is a rational B-spline surface. The edge attributes include an 11-dimensional one-hot vector $e_{\text{type}}\in\mathbb{R}^{11}$ for identifying edge types (e.g., circular, closed curve, elliptical, straight line, etc.), a vector $e_{\text{len}}\in\mathbb{R}^{1}$ to identify the type of the edge, and a 3-dimensional one-hot vector $e_{\text{conv}}\in\mathbb{R}^{3}$ to characterize the convexity of the edge (concave, convex, or smooth). The encoded information of attribute features is shown in Table 2 as indicated.
Table 2. Attribute Feature Encoding for Geometric Elements
Element Input Feature Encoder Layer Output
Face $f_{\text{type}}\in R^{9}$ Linear (9, 32) $h_{f,\text{type}}\in R^{32}$
$f_{\text{area}}\in R^{1}$ Linear (1, 32) $h_{f,\text{area}}\in R^{32}$
$f_{\text{cen}}\in R^{3}$ Linear (3, 32) $h_{f,\text{cen}}\in R^{32}$
$f_{\text{rat}}\in R^{1}$ Linear (1, 32) $h_{f,\text{rat}}\in R^{32}$
Edge $e_{\text{type}}\in R^{11}$ Linear (11, 64) $h_{e,\text{type}}\in R^{64}$
$e_{\text{len}}\in R^{1}$ Linear (1, 32) $h_{e,\text{len}}\in R^{32}$
$e_{\text{conv}}\in R^{3}$ Linear (3, 32) $h_{e,\text{conv}}\in R^{32}$
Finally, by concatenating the encoded features of the two types of geometric attributes and the geometric UV domain features, we obtained the complete geometric face and edge features, $H_{\text{face}}$ and $H_{\text{edge}}$ respectively:
(4) $H_{\text{face}}=\text{Concat}(h_{f,\text{geo}},h_{f,\text{type}},h_{f,\text{area}},h_{f,\text{cen}},h_{f,\text{rat}})\in\mathbb{R}^{256}$
(5) $H_{\text{edge}}=\text{Concat}(h_{e,\text{geo}},h_{e,\text{type}},h_{e,\text{len}},h_{e,\text{conv}})\in\mathbb{R}^{256}$
Figure 3. The entire process of constructing our dataset, (A) illustrates the selection of four components from the database, (B) shows the application of random rotation, translation, and duplication under geometric constraints to generate a new B-rep model, and (C) demonstrates the process of traversing and labeling all faces (L1, L2, L3, L4).
3.4. Transformer Block Module
Before feeding the encoded face features into our transformer architecture, we introduced a virtual face feature that is connected to all B-rep elements. This virtual face feature, within the transformer structure, can engage in deep interactions with the actual face features, thereby obtaining a global feature that represents the entire B-rep model.
This part consists of 8 layers of designed transformer modules, with each layer primarily incorporating Grouped Query Attention (GQA) (Ainslie et al., 2023), Root Mean Square Normalization(RMS Norm), and SwiGLU activation function. First, all the feature of the faces are represented as $H_{\text{face}}^{0}=\left[{h_{\text{face}}^{1}},{h_{\text{face}}^{2}},\ldots,{h_{\text{face}}^{N+1}}\right]\in\mathbb{R}^{{(N+1)}\times 256}$. Here, each $h_{\text{face}}^{i}$ represents the feature of a given face, $N+1$ indicates that it includes the features of the virtual face, and the superscript 0 in $H_{\text{face}}^{0}$ denotes the initial iteration of the overall feature representation.
First, all the input face features are fed into RMS Norm for normalization. Next, the normalized features are fed into the GQA for feature propagation. Following this, the features processed by the attention mechanism are fed into the Feed-Forward Network (FFN), and then combined with residual connections for further processing. The overall formula expressed as follows:
(6) $H_{\text{face}}^{t}{}^{\prime}=\text{Attention}\left(\text{RMS Norm}(H_{\text{face}}^{t-1})\right)+H_{\text{face}}^{t-1}$
(7) $H_{\text{face}}^{t}=\text{FFN}\left(\text{RMS Norm}(H_{\text{face}}^{t}{}^{\prime})\right)+H_{\text{face}}^{t}{}^{\prime}$
Our attention uses GQA, which divides the queries into multiple groups and independently computes attention within each group. It then concatenates the outputs of all groups $Q_{g}$ and passes them through a linear transformation to obtain the output, effectively reducing computational load and memory usage, and $t$ indicates the iteration times. The specific expression formula for GQA is as follows:
(8) $O=\left(\text{Concat}\left[\text{softmax}\left(\frac{Q_{g}K^{T}}{\sqrt{d_{k}}}+B_{\text{Att}}\right)V\right]_{g=1}^{G}\right)W_{o}$
where $O$ refers to the final output, $G$ indicates the upper limit of the number of groups. For each group $g$, the dot product of $Q_{g}$ and $K$ is first scaled by the square root of the key vector dimension $d_{k}$. This scaled product is then adjusted by the attention bias $B_{att}$, and the $softmax$ function is applied to obtain the attention weights. These weights are used to compute the weighted sum of the values $V$. The outputs of all groups are concatenated and passed through a linear transformation $W_{o}$ to produce the final output $O$.
In the FFN layer, we primarily use the SwiGLU activation function, which is mathematically expressed as follows:
(9) $\text{SwiGLU}=\text{Swish}(Wx+b)\otimes(Vx+c)$
here, $W$ and $V$ denote the learnable weight matrices that are applied to the input $x$, $b$ and $c$ are bias terms. The Swish activation function defined as $\text{Swish}(z)=z\cdot{sigmoid}(z)$.
Finally, based on the introduced global virtual face feature and the input encoded face features, the module outputs $f_{global}$ , representing the global feature of the B-rep model, and $f_{local}$ , representing the local feature.
3.5. Recognition Head
Based on the local feature representation of $f_{local}$ and the global feature $f_{global}$ output by the transformer module, we design a structure to fuse these two parts of features. First, we broadcast the global feature to the same dimension as the local feature and stack the two parts along a new dimension to obtain the integrated feature representation $F_{all}$. Then, we multiply the features $F_{all}$ by a learnable weight matrix $W_{w}$ and apply a $softmax$ activation function to obtain the weight representation of these features. Finally, we perform element-wise multiplication between this weight representation and the $F_{all}$ to generate the final output features. The specific formulas are as follows:
(10) $F_{all}=\text{$stack$}[f_{local},f_{global}]$
(11) $F_{\text{out}}=F_{\text{all}}\otimes\text{softmax}(F_{\text{all}}\cdot W_{w}+b_{w})$
where $stack$ denotes the operation of concatenation along a new dimension, $F_{all}$ represents the newly obtained features, $W_{w}$ represents the learnable weight matrix of the linear layer and $b_{w}$ denotes the bias term associated with it.
For the final feature recognition task in CAD models, we employed a MLP coupled with an $argmax$ activation function as the classifier to predict the geometric feature category $\hat{C}$ of each face. The mathematical expression is as follows:
(12) $\hat{C}=\arg\max(\text{MLP}(F_{\text{out}})\in\mathbb{R}^{N\times K})$
We choose the cross-entropy as our final loss function, which is as follow:
(13) $L=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{i,j}\log(\hat{C}_{i,j}+\epsilon)$
here, $\epsilon$ is a small positive constant added to the predicted probabilities to ensure numerical stability
4. Complex B-rep Feature (CBF) Dataset
We introduce a dataset named CBF to support research on complex geometric feature recognition, and describe the method for creating the corresponding dataset. The dataset comprises 20,000 CAD models in B-rep format. Each model is a B-rep model formed by boolean combinations of a base plate and three different geometric features on it. The faces of each geometric feature are labeled accordingly, with the label information stored in a separate JSON file to support research on model construction and training for these complex features.
For our data creation process, Figure 3 demonstrates the entire workflow. We first selected B-rep from public sources and separated the base plates and geometric features using relevant 3D modeling software. Subsequently, we applied random rotations, translations, and duplications to the three extracted geometric features on the base plate. During these operations, we ensured the rationality of the generated models by manipulating their geometric constraints. Specifically, for translation, the script moved the model along the normal and tangent vectors of the contact surface until it intersected with the base plate, ensuring proper attachment. Rotations were performed based on the geometric attributes of the contact surface (such as normal vectors and centroids) to avoid model distortion or detachment. Duplication was carried out via boolean operations to ensure that the newly generated components did not conflict with the base plate or other components. Finally, in the labeling stage, we traversed all faces of the composite model, determined which original component each face belonged to (base plate, feature A, feature B, feature C), and assigned a label to each face. This information was stored in a dictionary, with face indices as keys and corresponding component labels as values.
5. Experiments
5.1. Experimental Environment
We employed a single NVIDIA 4090 GPU and PyTorch-lightning V1.9.0 for network training, thereby emphasizing the lightweight characteristic of our network. In the training phase, AdamW was chosen as the optimizer, with an initial learning rate of 0.001. The parameters were set to $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon={1\times 10^{-8}}$ to guarantee training stability. Additionally, we employed the ReduceLROnPlateau (Al-Kababji et al., 2022) learning rate scheduling strategy, which dynamically adjusts the learning rate based on changes in validation loss. It is important to note that during the initial 5000 steps of training, we implemented a linear learning rate warm-up phase to help the model converge more effectively. The batch size was set to 64, and the entire training process lasted for a maximum of 200 epochs.
To evaluate the performance of our network, we used overall accuracy $A$, class accuracy $A_{c}$, and mean Intersection over Union (mIoU) to verify the performance of our model’s geometric feature recognition capabilities, mathematically expressed as follows:
(14) $A=\frac{F_{\text{c}}}{F_{\text{t}}}$
here, $F_{c}$ denotes the number of correctly classified B-rep faces, and $F_{t}$ is the total number of B-rep faces in the CAD model. Overall accuracy A calculates the proportion of correctly classified B-rep faces to the total number of B-rep faces in the CAD model.
(15) $A_{c}=\frac{1}{C}\sum_{c=1}^{C}\frac{\hat{y}_{c}=l_{c}}{y_{c}=l_{c}}$
here, $A_{c}$ denotes the average accuracy across all classes. $y_{c}$ represents the total number of B-rep faces in label $l_{c}$, while $\hat{y}_{c}$ indicates the number of correctly predicted B-rep faces in label $l_{c}$. The class accuracy $A_{c}$ represents the average accuracy across all geometric feature classes.
(16) $mIoU=\frac{1}{C}\sum_{c=1}^{C}\frac{\hat{y}_{c}=l_{c}\cap y_{c}=l_{c}}{\hat{y}_{c}=l_{c}\cup y_{c}=l_{c}}$
here, the parameter expressions in this part are the same as those described above for $A_{c}$. The mIoU is commonly used to evaluate the overlap between predicted results and true labels. In the context of our geometric feature recognition task, mIoU is calculated as the average ratio of correctly classified B-rep faces that share the same class labels in both the actual and predicted outputs.
5.2. Experimental Datasets
We conducted experiments on the MFInstSeg and MFTRCAD public datasets to evaluate the machining feature recognition ability of our model and compared it with mainstream deep learning methods. The results indicate that our model achieved state-of-the-art accuracy on these datasets. Additionally, we verified the complex feature recognition capability of our model on the proposed CBF dataset. The results indicate that our network also achieves state-of-the-art accuracy. For all datasets, the data were split into 70% for training, 15% for validation, and 15% for testing.
5.2.1. MFInstSeg Dataset
The MFInstSeg dataset comprises 62,495 CAD model files stored in B-rep format. It includes 24 different types of machining features, with each model containing 3 to 10 unique machining features. Table 3 shows the performance of our model and other mainstream models on this dataset.
Table 3. Recognition Performance on MFInstSeg Dataset
Network Accuracy(%) mIOU(%)
ASIN(Zhang et al., 2022) $86.46\pm 0.45$ $79.15\pm 0.82$
GATv2(Brody et al., 2021) $95.90\pm 0.20$ $93.03\pm 0.36$
GraphSAGE(Hamilton et al., 2017) $97.69\pm 0.06$ $95.70\pm 0.14$
GIN(Xu et al., 2018) $98.14\pm 0.03$ $96.52\pm 0.06$
DeeperGCN(Li et al., 2020) $99.03\pm 0.02$ $98.31\pm 0.01$
AAGNet(Wu et al., 2024) $99.15\pm 0.03$ $98.45\pm 0.04$
MFTRNet(Xia et al., 2024) $99.56\pm 0.02$ $98.43\pm 0.03$
BRepFormer(Ours) $\textbf{99.62}\pm\textbf{0.03}$ $\textbf{98.74}\pm\textbf{0.09}$
5.2.2. MFTRCAD Dataset
The MFTRCAD dataset comprises 28,661 CAD models stored in B-rep format. The authors further divided one of the traditional 24 machining feature categories into three subcategories, leading to a total of 26 distinct machining features in the dataset. Table 4 presents the performance of our model and other mainstream models on this dataset.
Table 4. Recognition Performance on MFTRCAD Dataset
Network Accuracy(%)
PointNet++(Qi et al., 2017) $67.89\pm 0.08$
DGCNN(Wang et al., 2019) $67.97\pm 0.07$
ASIN(Zhang et al., 2022) $68.57\pm 0.41$
Hierarchical CADNet(Colligan et al., 2022) $78.39\pm 0.03$
AAGNet(Wu et al., 2024) $79.45\pm 0.02$
MFTRNet(Xia et al., 2024) $89.88\pm 0.02$
BRepFormer(Ours) $\textbf{93.16}\pm\textbf{0.11}$
5.2.3. Complex Feature Dataset
In our CBF dataset, unlike previous datasets, this dataset requires the model to identify these three distinct geometric features along with the base plate. The experimental results of our model and other mainstream models on this dataset is presented in Table 5.
Table 5. Recognition Performance on CBF Dataset
Network Accuracy(%) Class Acc(%) mIoU(%)
AAGNet(Wu et al., 2024) 93.41 - 93.76
MFTRNet(Xia et al., 2024) 92.12 - 91.21
BRepFormer(Ours) 94.66 94.97 87.48
Although our network outperforms other comparative networks in terms of overall accuracy, it performs poorly in the mIoU metric. Analysis reveals that when the network identifies simple geometric features, the limited number of faces involved means that any misidentification significantly impacts overall accuracy. In contrast, when dealing with complex geometric features, the higher number of faces means that misidentification of some faces has a relatively limited impact on overall accuracy. Based on the above analysis, despite certain shortcomings in the mIoU metric, our network still maintains a leading position in overall recognition accuracy. This indicates that our network is more adept at recognizing complex, multi-faceted geometric features.
5.3. Ablation Study
In the ablation study, we focused on the impact of the input features of our model on its performance. We systematically removed these key features and tested the changes in model performance. All ablation experiments were conducted on our proposed CBF dataset.
5.3.1. Ablation Analysis of Geometric Features
The geometric features directly input into our BRepFormer  network include the UV domain geometric features and the attribute features of the B-rep. In this ablation study, a model with complete input features was used as the baseline, and then each of the three input features was removed one by one to generate the ablation models. The results shown in Table 6 indicate that the removal of any input feature leads to a decrease in feature recognition accuracy. Among them, the removal of attribute features causes the most significant performance drop, while the removal of UV domain geometric features has a less noticeable impact. This suggests that attribute features are more important than the input geometric features within our network architecture.
Table 6. Impact of Removing Different Geometric Features on BRepFormer  Performance
Input Accuracy(%) Class Acc(%) mIoU(%)
Full (baseline) 94.66 94.97 87.48
w/o Face Attr 92.34 (-2.32) 91.23 (-3.74) 84.10 (-3.38)
w/o Edge Attr 91.01 (-3.65) 90.01 (-4.96) 81.69 (-5.79)
w/o UV-grid 92.95 (-1.71) 91.85 (-3.12) 85.05 (-2.43)
5.3.2. Ablation Analysis of Topological Features
In our BRepFormer network, we focused on extracting four key topological features from CAD models to delineate the comprehensive relational structure of B-rep models. To demonstrate the effectiveness of the extracted features, we established a baseline model using the initial complete topological features and then conducted ablation experiments by progressively removing topological features on our CBF dataset. As shown in Table 7, the network’s accuracy decreased successively with the removal of each of the topological features. This indicates that the global topological matrices facilitate accurate understanding of complex 3D solid models by neural network.
Table 7. Impact of Removing Different Topological Features on BRepFormer  Performance
Input Accuracy(%) Class Acc(%) mIoU(%)
Full (baseline) 94.66 94.97 87.48
w/o $M_{d}$ 93.51 (-1.15) 92.73 (-2.24) 86.28 (-1.20)
w/o $M_{d}$, $M_{a}$ and $M_{c}$ 93.33 (-1.33) 92.29 (-2.68) 85.95 (-1.53)
w/o $M_{d}$, $M_{a}$, $M_{c}$ and $M_{e}$ 93.22 (-1.44) 92.17 (-2.80) 85.60 (-1.88)
5.4. Geometric Recognition Presentation
This section presents a visual demonstration of our network’s performance in geometric feature recognition. Figure 4 illustrates the network’s capability in identifying machining features, with the green highlights indicating the features that have been successfully recognized by the network. Figure 5, meanwhile, displays the outcomes of our network’s feature recognition on more complex geometric shapes.
Figure 4. Examples of recognized machining features (highlighted in green) in B-rep models
Figure 5. Examples of recognized complex features (highlighted in green) in B-rep models
6. Conclusion
In this paper, we introduce BRepFormer, a novel geometric feature recognition network based on the transformer architecture. Our network effectively extracts both geometric and topological information from CAD models, and incorporates an attention bias that integrates geometric and topological features to regulate information propagation within the transformer module. Furthermore, we propose the CBF dataset, which features more complex geometric and topological representations and is specifically designed for complex feature recognition tasks. Finally, BRepFormer  achieves state-of-the-art accuracy on the public MFInstSeg and MFTRCAD datasets, as well as our CBF dataset, thereby demonstrating its superiority in both machining feature recognition and complex geometric feature recognition tasks.